OpenStudy (anonymous):

What is the rate of change of: 5x + 3y = -2. How did you find the rate of change?

OpenStudy (anonymous):

rate of change wiith respect to what?

OpenStudy (across):

assuming y(x) rearrange 3y=-2-5x divide y=-5/3x-2/3 now you tell me